# Design of Bayesian A/B Tests Controlling False Discovery Rates and Power

### v3

@@ 7 Discussion
In this paper, we put forward an economical framework to design Bayesian A/B tests involving multiple metrics while satisfying criteria for the FDR and average power. This framework recommends optimal sample sizes and decision thresholds that are tailored to the specified dependence structure between the hypothesis tests and accommodates uncertainty regarding which hypotheses are false. The efficiency of this framework stems from considering a proxy for the joint sampling distribution of posterior probabilities based on large-sample theory to motivate estimating the true sampling distributions at only two sample sizes. This approach significantly lessens the computational overhead required to design Bayesian A/B tests, making them much more attractive and accessible to practitioners that want to control the FDR. The posterior probabilities we compute can also be repurposed to (i) construct bootstrap confidence intervals that account for the impact of simulation variability and (ii) investigate the suitability of various sample sizes and decision thresholds using contour plots.
The methods proposed in this paper could be extended in various aspects to accommodate more complex designs, in the context of A/B testing or otherwise. Our work in this article enforced the constraint that the decision threshold $\gamma$ is the same for all hypotheses. For certain groups of A/B tests, we may want to allow different decision thresholds $\{\gamma_{k}\}_{k=1}^{K}$ for each of the $K$ hypotheses. Theorem 1 of this paper could be applied in that context, but further work is required to develop a general method to determine an optimal set of decision thresholds that controls the FDR.
When considering A/B tests with multiple metrics, it may be useful to consider design criteria other than the FDR or the FWER. This consideration could be framed using the three types of metrics defined by Kohavi et al. (2020) based on the evidence supporting metric improvement or deterioration in group B. In particular, positive and negative metrics respectively have statistically significant evidence of improvement and deterioration, and flat metrics have no statistically significant evidence of improvement or deterioration. Kohavi et al. (2020) recommended implementing a new online experience if all metrics are flat or positive, with at least one metric being positive. To classify the $k^{\text{th}}$ metric as positive, negative, or flat, we would need to compare $Pr(H_{1,k}~{}|~{}data)$ to both lower and upper decision thresholds. We could develop design methods that control the probabilities of incorrectly and correctly implementing new online experiences in this more complex framework.

### v4

@@ 6 Discussion
In this paper, we put forward an economical framework to design Bayesian A/B tests involving multiple metrics while satisfying criteria for the FDR and average power. This framework recommends optimal sample sizes and decision thresholds that are tailored to the specified dependence structure between the hypothesis tests and accommodates uncertainty regarding which hypotheses are false. The efficiency of this framework stems from considering a proxy for the joint sampling distribution of posterior probabilities based on large-sample theory to motivate estimating the true sampling distributions at only two sample sizes. This approach significantly decreases the computational overhead required to design Bayesian A/B tests, making them much more attractive and accessible to practitioners that want to control the FDR.
The methods proposed in this paper could be extended in various aspects. Further work could develop more sophisticated and efficient optimization schemes to determine decision thresholds $\{\gamma_{k}\}_{k=1}^{K}$ that control the FDR. It may also be useful to consider design criteria for A/B tests other than the FDR or the FWER. This consideration could be framed using the three types of metrics defined by Kohavi et al. (2020) based on the evidence supporting metric improvement or deterioration in group B. In particular, positive and negative metrics respectively have statistically significant evidence of improvement and deterioration, and flat metrics have no statistically significant evidence of improvement or deterioration. Kohavi et al. (2020) recommended implementing a new online experience if all metrics are flat or positive, with at least one metric being positive. To classify the $k^{\text{th}}$ metric as positive, negative, or flat, we would need to compare $Pr(H_{1,k}~{}|~{}data)$ to both lower and upper decision thresholds in an alternative design framework.